Answer it in 500 words with introduction, concept(body), conclusion

Define elasticity of supply and find the price from the given statement: If Es of a good is 2 and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units.


1
Expert's answer
2021-05-14T10:56:06-0400

Price elasticity of supply measures the responsiveness to the supply of a good or service after a change in its market price. Elastic means the product is considered sensitive to price changes. Inelastic means the product is not sensitive to price movements.

"Ed = \frac{250 - 200} {P2 - 8} ×\frac{P2 + 8} {250+200} = 2,"

18P2 - 144 = P2 + 8,

17P2 = 152,

P2 = 8.94.